OpenStudy (anonymous):

-3(2n+2)=9 solve for n.

OpenStudy (solomonzelman):

divide both sides by -3 subtract 2 from both sides divide both sides by 2

OpenStudy (anonymous):

got -2.5?

OpenStudy (solomonzelman):

yup!

OpenStudy (anonymous):

I have to write my answer as a simplified fraction or integer

OpenStudy (solomonzelman):

simplified fraction, then -5/2. mixed number -2, 1/2 you can't write this as an integer though, cuz it is not a whole number.

OpenStudy (anonymous):

Can you help with another question?

OpenStudy (solomonzelman):

Sure.

OpenStudy (anonymous):

4(5z-1)-2(2+7)=5(z+1) z = ?

OpenStudy (solomonzelman):

4(5z-1)-2(2+7)=5(z+1) expand each 4(5z-1)=? 2(2+7)=? 5(z+1)=?

OpenStudy (anonymous):

each variable equals one? so its 4(5(1)-1) 2(2+7) 5(1+1) ??

OpenStudy (solomonzelman):

don't substitute for z, just expand each of those like I asked.

OpenStudy (anonymous):

20z - 4 ? 18 ? 10 ?

OpenStudy (solomonzelman):

5(z+1) is 5z-5 so 20z - 4 +18=5z-5 add like terms.

OpenStudy (solomonzelman):

My bad 20z - 4 +18=5z+5

OpenStudy (solomonzelman):

15z+14=5 15z=5-14 15z=-9 z=-3/5